# Task-Oriented Communications for Visual Navigation with Edge-Aerial Collaboration in Low Altitude Economy

### v3

@@ Abstract
To support the Low Altitude Economy (LAE), precise unmanned aerial vehicles (UAVs) localization in urban areas where global positioning system (GPS) signals are unavailable. Vision-based methods offer a viable alternative but face severe bandwidth, memory and processing constraints on lightweight UAVs. Inspired by mammalian spatial cognition, we propose a task-oriented communication framework, where UAVs equipped with multi-camera systems extract compact multi-view features and offload localization tasks to edge servers. We introduce the Orthogonally-constrained Variational Information Bottleneck encoder (O-VIB), which incorporates automatic relevance determination (ARD) to prune non-informative features while enforcing orthogonality to minimize redundancy. This enables efficient and accurate localization with minimal transmission cost. Extensive evaluation on a dedicated LAE UAV dataset shows that O-VIB achieves high-precision localization under stringent bandwidth budgets. Code and dataset will be made publicly available: github.com/fangzr/TOC-Edge-Aerial.
@@ I Introduction
  • We propose O-VIB, a novel encoder incorporating orthogonality constraints and ARD-driven sparsification, enabling compact feature encoding that supports accurate localization at low communication cost.
@@ III-B2 Orthogonality Under the IB Objective
In our VIB-based encoder design, we compress the multi-view feature $\mathbf{x}\in\mathbb{R}^{M\times d}$ into a low-dimensional latent representation $\mathbf{z}\in\mathbb{R}^{k}$. Following the IB principle, we maintain a small mutual information $\beta I(\mathbf{x};\mathbf{z})$ to limit bandwidth usage while retaining sufficient information about $\mathbf{x}$ for accurate prediction of $\mathbf{y}$. To optimize the utilization of the limited information budget, we impose approximate row-orthogonality on the encoder’s weight matrix $\mathbf{W}$. The proposition 1 shows that if $\mathbf{W}\mathbf{W}^{\top}$ is close to an identity matrix, the variance of each latent coordinate remains close to the average $\frac{1}{k}\operatorname{Tr}(\mathbf{W}\Sigma_{x}\mathbf{W}^{\top})$. Consequently, no latent dimension collapses to near-zero variance, thereby avoiding redundancy in the latent representation.
@@ IV Performance Evaluation
As shown in Fig. 5, we collect a new dataset for visual navigation of UAVs in the CARLA simulator that mimics GNSS-denied flight over eight representative maps in cities [^1] The code, demos and dataset will be made publicly available at: github.com/fangzr/TOC-Edge-Aerial. A UAV flies at a constant height following waypoints aligned with the road, changing direction randomly. Five onboard cameras capture images from different angles and directions, recording RGB, semantic, and depth images at a 400 × 300 pixel resolution. A total of 357,690 multi-view frames are recorded, each labeled with the precise localization and rotation. As shown in Fig. 5, the framework is deployed on real hardware devices to evaluate the computation and communication latency of our proposed methods. Each UAV carries a Jetson Orin NX 8GB that encodes five camera streams and transmits compressed features to the nearby roadside units (RSUs) through wireless channels (IEEE 802.11). Two classes of RSU are deployed: (i) Relay RSU: Raspberry Pi 5 16 GB that forward data by Gigabit Ethernet to a cloud edge server when overloaded; (ii) Edge RSU: Jetson Orin NX Super 16 GB that performs on-board inference.
@@
Fig. 6 shows localization accuracy as the UAV–RSU link budget (in KB/s) is varied. Each scheme’s parameter set (e.g. latent dimension for VIB; quality factor for JPEG/H.264/H.265/WebP) is swept to identify Pareto‐optimal points—minimizing error per KB/s. Features are encoded onboard, sent over IEEE 802.11, decoded at the RSU, and matched to a geo-tagged database; the nearest neighbor index yields the pose estimate. When network throughputs are above 50 KB/s, all methods converge to a mean error of 10 m. When the bottleneck falls below 10 KB/s, O-VIB degrades most gracefully: at 8 KB/s it still achieves less than 10 m error, representing a 42.1% reduction versus vanilla VIB and a 62.6% reduction versus WebP. Embedding orthogonality thus prunes redundant latent dimensions while preserving task-critical information, making O-VIB far more robust under severe bandwidth constraints.

### v4

@@ Abstract
To support the Low Altitude Economy (LAE), it is essential to achieve precise localization of unmanned aerial vehicles (UAVs) in urban areas where global positioning system (GPS) signals are unavailable. Vision-based methods offer a viable alternative but face severe bandwidth, memory and processing constraints on lightweight UAVs. Inspired by mammalian spatial cognition, we propose a task-oriented communication framework, where UAVs equipped with multi-camera systems extract compact multi-view features and offload localization tasks to edge servers. We introduce the Orthogonally-constrained Variational Information Bottleneck encoder (O-VIB), which incorporates automatic relevance determination (ARD) to prune non-informative features while enforcing orthogonality to minimize redundancy. This enables efficient and accurate localization with minimal transmission cost. Extensive evaluation on a dedicated LAE UAV dataset shows that O-VIB achieves high-precision localization under stringent bandwidth budgets. Code and dataset will be made publicly available at: github.com/fangzr/TOC-Edge-Aerial.
@@ I Introduction
  • We propose O-VIB, a novel encoder incorporating orthogonality constraints and ARD-driven sparsification, that enables compact feature encoding that supports accurate localization at low communication cost.
@@ III-B2 Orthogonality Under the IB Objective
In our VIB-based encoder design, we compress the multi-view feature $\mathbf{x}\in\mathbb{R}^{M\times d}$ into a low-dimensional latent representation $\mathbf{z}\in\mathbb{R}^{k}$. Following the IB principle, we aim to minimize the mutual information $I(\mathbf{x};\mathbf{z})$ (weighted by $\beta$) to limit bandwidth usage while ensuring $\mathbf{z}$ retains sufficient information about $\mathbf{x}$ for accurately predicting $\mathbf{y}$. To optimize the utilization of the limited information budget, we impose approximate row-orthogonality on the encoder’s weight matrix $\mathbf{W}$. The proposition 1 shows that if $\mathbf{W}\mathbf{W}^{\top}$ is close to an identity matrix, the variance of each latent coordinate remains close to the average $\frac{1}{k}\operatorname{Tr}(\mathbf{W}\Sigma_{x}\mathbf{W}^{\top})$. Consequently, no latent dimension collapses to near-zero variance, thereby avoiding redundancy in the latent representation.
@@ IV Performance Evaluation
As shown in Fig. 5, we have collected a new dataset for visual navigation of UAVs in the CARLA simulator that mimics GNSS-denied flight over eight representative maps in cities [^1] The code, demos and dataset will be made publicly available at: github.com/fangzr/TOC-Edge-Aerial. A UAV flies at a constant height following waypoints aligned with the road, changing direction randomly. Five onboard cameras capture images from different angles and directions, recording RGB, semantic, and depth images at a 400 × 300 pixels. A total of 357,690 multi-view frames are recorded, each labeled with the precise localization and rotation. As shown in Fig. 5, the framework is deployed on real hardware devices to evaluate the computation and communication latency of our proposed methods. Each UAV carries a Jetson Orin NX 8GB that encodes five camera streams and transmits compressed features to the nearby roadside units (RSUs) through wireless channels (IEEE 802.11). Two classes of RSU are deployed: (i) Relay RSU: Raspberry Pi 5 16 GB that forwards data by Gigabit Ethernet to a cloud edge server when overloaded; (ii) Edge RSU: Jetson Orin NX Super 16 GB that performs on-board inference.
@@
Fig. 6 shows localization accuracy as the UAV–RSU link budget (in KB/s) is varied. Each scheme’s parameter set (e.g. latent dimension for VIB; quality factor for JPEG/H.264/H.265/WebP) is swept to identify Pareto‐optimal points—minimizing error per KB/s. Features are encoded onboard, sent over IEEE 802.11, decoded at the RSU, and matched to a geo-tagged database; the nearest neighbor index yields the pose estimate. When network throughput is above 50 KB/s, all methods converge to a mean error of 10 m. When the bottleneck falls below 10 KB/s, O-VIB degrades most gracefully: at 8 KB/s it still achieves less than 10 m error, representing a 42.1% reduction versus vanilla VIB and a 62.6% reduction versus WebP. Embedding orthogonality thus prunes redundant latent dimensions while preserving task-critical information, making O-VIB far more robust under severe bandwidth constraints.